a 10g bullet having a charge of 4.00 micro C is fired at aspeed of 270m/s in a horizontal direction.avertical magnetic field of 500 micro T exists in the space.find the deflection of the bullet due to the magnetic field  as it travels through 100m.

Grade:12

1 Answers

askIITianexpert IITDelhi
14 years ago

F=qvB=ma   ;    a=qvB/m=54micro units(mtrs./sec.2)

S=1/2at2  where t=100/270     ;  S=100/27 micro mtrs.  :this is the deflection.